# LLM-Barber: Block-Aware Rebuilder for Sparsity Mask in One-Shot for Large Language Models

### v1

@@ Introduction
Table 1: WikiText-2 perplexity comparison for pruning LLM models at 50% sparsity rate. Bold results show improvements of integrating LLM-Barber. Underscored results indicate best performance in each LLM. WR represents weight reconstruction.
Output: Rebuilt sparsity masks $\{\mathbf{M_{r}}^{l}\}^{L}_{l=1}$.
1:  $\mathbf{E_{i}}\leftarrow\text{Block}\mathbf{(W,X)}-\text{Block}\mathbf{(W\odot M_{i},X)}$
2:  $\{\mathbf{G}^{l}\}^{L}_{l=1}\leftarrow$ backpropagation for gradients via $\mathbf{E_{i}}$.
3:  for $l$ in $\{1,2,\dots,L\}$ do
4:     $\mathbf{M_{r}}^{l}\leftarrow\mathbf{M_{i}}^{l}$triangleright Initialize rebuilt sparsity mask.
5:     $\mathbf{S}^{l}\leftarrow|\mathbf{W}^{l}|\cdot|\mathbf{G}^{l}|$triangleright Obtain importance score.
[…]
7:     for $n$ in $\{1,2,\dots,N\}$ do
8:        Obtain growing index $gi,gj$ via Eq.(10)
9:        Obtain pruning index $pi,pj$ via Eq.(11)
10:        $\mathbf{M_{r}}^{l}_{pi,pj}=1$triangleright Weight Growing.
11:        $\mathbf{M_{r}}^{l}_{gi,gj}=0$triangleright Weight Pruning.
12:     end for
13:  end for
14:  $\mathbf{E_{r}}\leftarrow\text{Block}\mathbf{(W,X)}-\text{Block}\mathbf{(W\odot M_{r},X)}$
15:  Identify improvement by comparing error $\mathbf{E_{i}}$ and $\mathbf{E_{r}}$.
16:  return rebuilt sparsity masks $\{\mathbf{M_{r}}^{l}\}^{L}_{l=1}$.
Algorithm 1 Pseudocode of LLM-Barber.
Procedure
Like most post-training pruning methods (Han et al. 2015; Sun et al. 2023), LLM-Barber is executed within a single global LLM forward pass, with local backward passes for gradient computation in each block. Figure 2 illustrates the LLM-Barber workflow, which consists of four stages:
  • Sparsity Mask Initialization. We apply a post-training pruning technique to initialize a preliminary sparsity mask from the dense network.
  • Block-aware Reconstruction Error Computation. We use a block-aware reconstruction error to evaluate the discrepancy between the dense and sparse model outputs.
  • Back-propagation for Gradients. Gradients are automatically derived via back-propagation, and the product of weights and gradients serves as the pruning metric.
  • Sparsity Mask Rebuilding. Masks are sorted based on the pruning metric, unpruned weights in ascending order while pruned weights in descending order. We rebuild the weight masks by growing newly significant weights and pruning those became non-salient with a certain percentage named mask rebuilding ratio.
Structured N:M Sparsity
While LLM-Barber primarily targets unstructured sparsity, it can be adapted for structured N:M sparsity. Here, groups of M weights are pruned to retain only N non-zero weights. During mask rebuilding, LLM-Barber divides each M-group into N pairs (each pair contains one pruned and one non-pruned weight), then sorts these pairs by output channel to determine which are mask rebuilding pairs. This method creates an optimal sparse mask, leveraging N:M sparsity while maintaining model performance.
Experiment
Experiment Settings
@@
Models & Datasets. LLM-Barber is evaluated on the LLaMA family, including LLaMA-7B/13B (Touvron et al. 2023a), LLaMA2-7B/13B (Touvron et al. 2023b), and LLaMA3-8B (AI@Meta 2024), as well as the OPT model series: OPT-6.7B/13B (Zhang et al. 2022). Notably, LLM-Barber is broadly applicable to any Transformer-based LLMs with Self-Attention and MLP blocks. Following previous works, we use 128 segments of 2048 tokens from the C4 dataset (Raffel et al. 2020) for mask rebuilding.
Evaluation. To comprehensively assess LLM-Barber, we conduct rigorous evaluations on perplexity and zero-shot accuracy. Perplexity is measured on the validation sets of benchmarks such as WikiText-2 (Merity et al. 2017), PTB (Marcus et al. 1994), and C4 (Raffel et al. 2020). Zero-shot accuracy is assessed using the EleutherAI LM Harness (Gao et al. 2023) across six benchmarks: BoolQ (Clark et al. 2019), RTE (Wang et al. 2018), HellaSwag (Zellers et al. 2019), ARC Easy and Challenge (Clark et al. 2018), and OpenbookQA (Mihaylov et al. 2018).
Baselines. The results of LLM-Barber are compared with the following established post-training pruning methods:
  • Magnitude Pruning (Han et al. 2015) eliminates weights based only on their magnitudes;
  • SparseGPT (Frantar and Alistarh 2023) identifies weights importance by using second-order information;
  • Wanda (Sun et al. 2023) determines weights to be pruned by the weight magnitude multiplied by input activation.
Main Results
Unstructured Sparsity. LLM-Barber effectively prunes the LLaMA and OPT models, achieving 50% unstructured sparsity without requiring supplementary weight reconstruction, as detailed in Table 1. LLM-Barber demonstrates the capability to rebuild the sparsity masks initialized by other pruning methods in a single forward pass, significantly outperforming conventional pruning baselines. In the LLaMA3-8B model, LLM-Barber creates a new sparsity mask that reduces perplexity to 9.451, a substantial improvement over the Wanda baseline of 9.821. Notably, LLM-Barber achieves robust improvements even with poorly performing initial sparsity masks, such as magnitude pruning, where it reduces perplexity from 205.5 to 10.98, which is an impressive and substantial enhancement.
@@
w/ LLM-Barber 22.04 1.07e2 5.70e2 7.07e3
Table 2: WikiText-2 perplexity performance for pruning LLaMA3-8B at varying sparsity rate.
Varying Sparsity Levels. We conduct experiments on varying sparsity levels for unstructured pruning in LLaMA3-8B as shown in Table 2. LLM-Barber consistently increases perplexity across all initialization methods, with magnitude pruning showing the most significant improvement.
@@
w/ LLM-Barber 2:4 10.14 8.517 9.806
Table 3: WikiText-2 perplexity comparison for pruning LLaMA family with structured N:M pattern.
Structured N:M Sparsity. In contrast to unstructured sparsity, employing N:M fine-grained sparsity can provide more tangible acceleration benefits when leveraging NVIDIA Ampere’s sparse tensor cores (Choquette et al. 2021).
@@
Table 4: Zero-shot performance comparison for pruning LLaMA series on six evaluated tasks at 50% sparsity rate. Bold results show improvements of integrating LLM-Barber. Underscored results indicate the best performance in each zero-shot tasks.
Therefore, we also evaluate the effectiveness of our LLM-Barber in partial LLaMA models on the N:M fine-grained sparsity pattern as shown in Table 3.
Zero-shot Performance. Following previous works (Frantar and Alistarh 2023; Sun et al. 2023), we evaluated the LLaMA models on six diverse zero-shot tasks. The results are summarized in Table 4, where models are pruned to unstructured 50% sparsity. Averaging the accuracy across the six evaluated tasks, it becomes apparent that LLM-Barber possesses the capability to identify a more effective network than those obtained via the initialization methods. For tasks such as BoolQ, RTE, and ARC-e, LLM-Barber consistently outperforms the baseline pruning techniques across the entire LLaMA model suite. However, it is worth noting that there is no single universally superior performer for the remaining tasks in the evaluation set, with the initial pruning methods sometimes matching or even marginally exceeding the results obtained through LLM-Barber.
Mask Rebuilding Ratio Selection
Figure 3: Relationship between importance score distribution of mask rebuilding pair and WikiText-2 perplexity results at varying pruning granularities in LLaMA-7B. The green dashed line indicates the best choice of mask rebuilding ratio.
@@
We have plotted the score distribution of mask rebuilding pairs for various pruning granularity and initialization methods on LLaMA-7B, as well as perplexity performance corresponding to mask rebuilding ratios in Figure 3. The results reveal a strong correlation between the distribution of outliers in mask rebuilding pairs and the optimal mask rebuilding ratio. For instance, outliers are significantly distributed within the top 10% with Magnitude pruning, thus selecting a 10% mask rebuilding ratio yields the overall optimal solution. Similarly, with the initialization method Wanda, a significant outlier distribution around the 1% mark corresponds to optimal results near a 1% mask rebuilding ratio. Thus, LLM-Barber can preemptively narrow the search range for the optimal mask rebuilding ratio by analyzing outlier distributions. This approach allows LLM-Barber to adapt flexibly to various reconstruction masks without extensive searching.
@@
Ablation Study
Given the outstanding potential of LLM-Barber, we have comprehensively analyzed the impact of three critical factors to assess the robustness and effectiveness of our LLM-Barber method: pruning granularity, pruning metric, and calibration data size. These factors were chosen to gain a deeper understanding of how different configurations influence the performance of pruned models and to demonstrate LLM-Barber’s versatility across various settings.
Pruning Granularity. LLM-Barber dynamically selects different pruning granularities to adapt to varying initialization methods. In this paper, we evaluated the impact of four levels of granularities: block-wise, layer-wise, input-wise, and output-wise pruning, as shown in Table 5. For magnitude pruning, the block-wise granularity yields the best performance, while the output-wise granularity delivers the lowest perplexity for Wanda pruning. By tailoring the pruning granularity to the particular pruning approach, LLM-Barber can consistently achieve optimal model performance.
@@
w/ LLM-Barber Block Layer Input Output
Magnitude (205.5) 10.98 11.06 72.48 11.81
SparseGPT (9.994) 9.380 9.348 9.418 9.567
Wanda (9.821) 9.626 9.633 9.849 9.451
Table 5: Ablation of pruning granularity in LLaMA3-8B. Perplexity in brackets show the WikiText-2 score without LLM-Barber. Bold results show best granularity of this row.
Pruning Metric. We analyze the effect of different pruning metrics, focusing on weight magnitude, gradient magnitude, and the product of weight and gradient. As seen in Table 6, the product of weight and gradient consistently outperforms other metrics, achieving the lowest perplexity of 10.98 for Magnitude, 9.348 for SparseGPT, and 9.451 for Wanda. This confirms the effectiveness of our product-based pruning metric across various methods.
@@
w/ LLM-Barber $|\mathbf{W}|$ $|\mathbf{\partial E/\partial W}|$ $|\mathbf{W}|\cdot|\mathbf{\partial E/\partial W}|$
Magnitude (205.5) 186.5 14.43 10.98
SparseGPT (9.994) 9.544 9.457 9.348
Wanda (9.821) 9.880 9.699 9.451
Table 6: Ablation of pruning metric in LLaMA3-8B. Bold results support the pruning metric $|\mathbf{W}|\cdot|\mathbf{\partial E/\partial W}|$.
@@
Figure 4: Ablation of calibration data size in LLaMA3-8B. LLM-Barber is robust across varying calibration data size.
Conclusion
We propose LLM-Barber (Block-Aware Rebuilder for Sparsity Mask in One-Shot), a novel framework that rebuilds sparsity mask to optimize LLM post-training pruning. By integrating a block-aware approach across Self-Attention and MLP blocks, LLM-Barber effectively reallocates sparsity to improve accuracy without the need for extensive fine-tuning. Specifically, LLM-Barber identifies novel importance score after mask initialization and rebuilds the sparsity mask with mask rebuilding pairs, simultaneously applying new sparsity masks to weights that have become less critical, thereby optimizing overall model performance. By utilizing the novel pruning metric as the product of weights and gradients, our approach enables precise and efficient reallocation of sparsity mask. Extensive experiments on pruning LLaMA series models demonstrate that LLM-Barber achieves state-of-the-art results in both perplexity and zero-shot performance within the domain of post-training pruning.

### v2

@@ III-D Structured N:M Sparsity
IV Experiment
IV-A Experiment Settings
@@
Models & Datasets. LLM-Barber is evaluated on the LLaMA family, including LLaMA-7B/13B [16], LLaMA2-7B/13B [17], and LLaMA3-8B [18], as well as the OPT model series: OPT-6.7B/13B [19]. Notably, LLM-Barber is broadly applicable to any Transformer-based LLMs with Self-Attention and MLP blocks. Following previous works, we use 128 segments of 2048 tokens from the C4 dataset [20] for mask rebuilding.
Evaluation. To comprehensively assess LLM-Barber, we conduct rigorous evaluations on perplexity and zero-shot accuracy. Perplexity is measured on the validation sets of benchmarks such as WikiText-2 [21], PTB [22], and C4 [20]. Zero-shot accuracy is assessed using the EleutherAI LM Harness [23] across six benchmarks: BoolQ [24], RTE [25], HellaSwag [26], ARC Easy and Challenge [27], and OpenbookQA [28].
Baselines. The results of LLM-Barber are compared with the following established post-training pruning methods: (1) Magnitude Pruning [15] eliminates weights based only on their magnitudes; (2) SparseGPT [4] identifies weights importance by using second-order information; (3) Wanda [8] determines weights to be pruned by the weight magnitude multiplied by input activation.
IV-B Main Results
Unstructured Sparsity. LLM-Barber effectively prunes the LLaMA and OPT models, achieving 50% unstructured sparsity without requiring supplementary weight reconstruction, as detailed in Table I. LLM-Barber demonstrates the capability to rebuild the sparsity masks initialized by other pruning methods in a single forward pass, significantly outperforming conventional pruning baselines. In the LLaMA3-8B model, LLM-Barber creates a new sparsity mask that reduces perplexity to 9.451, a substantial improvement over the Wanda baseline of 9.821. Notably, LLM-Barber achieves robust improvements even with poorly performing initial sparsity masks, such as magnitude pruning, where it reduces perplexity from 205.5 to 10.98, which is an impressive and substantial enhancement.
Varying Sparsity Levels. We conduct experiments on varying sparsity levels for unstructured pruning in LLaMA3-8B as shown in Table II. LLM-Barber consistently increases perplexity across all initialization methods, with magnitude pruning showing the most significant improvement.
TABLE II: WikiText-2 perplexity performance for pruning LLaMA3-8B at varying sparsity rate.
@@ IV-B Main Results
w/ Barber 22.04 1.07e2 5.70e2 7.07e3
Structured N:M Sparsity. In contrast to unstructured sparsity, employing N:M fine-grained sparsity can provide more tangible acceleration benefits when leveraging NVIDIA Ampere’s sparse tensor cores [29]. Therefore, we also evaluate the effectiveness of our LLM-Barber in partial LLaMA models on the N:M fine-grained sparsity pattern as shown in Table III.
TABLE III: WikiText-2 perplexity comparison for pruning LLaMA family with structured N:M pattern.
@@
w/ Barber 2:4 10.14 8.517 9.806
TABLE IV: Zero-shot performance comparison of LLaMA series on six tasks at 50% sparsity rate. Bold results show improvements of integrating LLM-Barber. Underscored results indicate the best performance for each tasks.
@@
Zero-shot Performance. Following previous works [4, 8], we evaluated the LLaMA models on six diverse zero-shot tasks. The results are summarized in Table IV, where models are pruned to unstructured 50% sparsity. Averaging the accuracy across the six evaluated tasks, it becomes apparent that LLM-Barber possesses the capability to identify a more effective network than those obtained via the initialization methods. For tasks such as BoolQ, RTE, and ARC-e, LLM-Barber consistently outperforms the baseline pruning techniques across the entire LLaMA model suite. However, it is worth noting that there is no single universally superior performer for the remaining tasks in the evaluation set, with the initial pruning methods sometimes matching or even marginally exceeding the results obtained through LLM-Barber.
TABLE V: Comparison of DSnoT vs. LLM-Barber on LLaMA3-8B at 50% unstructured sparsity. We report WikiText-2 perplexity (lower is better) and zero-shot accuracy (%) on five downstream tasks. Bold indicates best result for each setting.
Method WikiText-2 ARC-E ARC-C BoolQ OBQA RTE
Dense (no pruning) 6.136 75.29 41.89 75.08 34.40 66.79
Magnitude (50%) 205.5 46.59 25.09 42.87 22.00 53.07
    w/ DSnoT 47.52 60.52 32.16 49.66 25.20 54.87
    w/ LLM-Barber 10.98 72.05 37.49 72.72 27.40 54.87
Wanda (50%) 9.821 71.38 40.10 78.41 29.40 60.29
    w/ DSnoT 9.562 71.25 40.01 75.99 29.00 61.81
    w/ LLM-Barber 9.451 71.46 39.85 78.59 29.00 61.37
Comparison with DSnoT. We further compare LLM-Barber with the recent training-free pruning method DSnoT [7] to highlight their differences in approach and performance. DSnoT also aims to improve sparse LLMs without weight retraining, but it does so by iteratively pruning and regrowing weights in multiple cycles to minimize reconstruction In contrast, LLM-Barber performs all mask rebuilding in a single global forward pass (with localized backward gradient computations per block), making it a true one-shot method. This fundamental distinction gives LLM-Barber an efficiency edge: it avoids the repeated passes of DSnoT, enabling faster one-shot deployment. Moreover, DSnoT imposes a fixed number of iterations (a maximum cycle count) for mask update in each layer, which can lead to sub-optimal results when the initial mask is poor. LLM-Barber is not bound by a preset iteration limit; instead, it dynamically rebuilds the mask based on our weight-importance metric until all significant weights are retained. This adaptability allows LLM-Barber to recover much more performance, especially starting from an inferior initial mask. Importantly, LLM-Barber’s block-aware strategy provides greater flexibility via global coordination across Transformer blocks, whereas DSnoT treats layers more independently (resembling layerwise fine-tuning). By coordinating mask updates at the block level, LLM-Barber better preserves end-to-end model quality.
In summary, qualitatively LLM-Barber offers higher efficiency, flexibility, and overall performance than DSnoT. Its one-shot block-aware mask rebuilding avoids DSnoT’s iterative overhead and limitation, yielding a more globally coordinated pruning outcome. These advantages make LLM-Barber well-suited for scenarios that require rapid one-shot compression for LLM deployment without sacrificing accuracy.
[…]
IV-C Mask Rebuilding Ratio Selection
Figure 3: The importance score distribution of mask rebuilding pairs and WikiText-2 perplexity results at varying pruning granularities in LLaMA-7B, with the green dashed line marking the optimal mask rebuilding ratio.
@@ IV-C Mask Rebuilding Ratio Selection
We plotted the score distribution of mask rebuilding pairs across various pruning granularity and initialization methods on LLaMA-7B, along with perplexity performance corresponding to mask rebuilding ratios (Figure 3). The results reveal a strong correlation between the distribution of outliers in mask rebuilding pairs and the optimal mask rebuilding ratio. For instance, outliers are significantly distributed within the top 10% with Magnitude pruning, thus selecting a 10% mask rebuilding ratio yields the overall optimal solution. Similarly, the Wanda initialization method shows a notable outlier distribution around the 1% mark corresponds to optimal results near a 1% mask rebuilding ratio. Thus, LLM-Barber can preemptively narrow the search range for the optimal mask rebuilding ratio by analyzing outlier distributions, allowing for flexible adaptation to various reconstruction masks without extensive searching.
@@
TABLE VI: Pruning granularity ablation in LLaMA3-8B. Bold results show best granularity of each row.
@@
w/ Barber Block Layer Input Output
Magnitude 10.98 11.06 72.48 11.81
SparseGPT 9.380 9.348 9.418 9.567
Wanda 9.626 9.633 9.849 9.451
IV-D Ablation Study
Given the significant potential of LLM-Barber, we analyzed three critical factors to assess the robustness and effectiveness of our method: pruning granularity, pruning metric, and calibration data size. These factors were chosen to gain a deeper understanding of how different configurations affect the performance of pruned models and to demonstrate LLM-Barber’s versatility across various settings.
TABLE VII: Ablation of pruning metric in LLaMA3-8B. Bold results support the pruning metric $|\mathbf{W}||\mathbf{\partial E/\partial W}|$.
@@ IV-D Ablation Study
w/ Barber $||\mathbf{W}|$$||\mathbf{\partial E/\partial W}|$$|\mathbf{W}||\mathbf{\partial E/\partial W}|$
Magnitude 186.5 14.43 10.98
SparseGPT 9.544 9.457 9.348
Wanda 9.880 9.699 9.451
Pruning Granularity. LLM-Barber dynamically selects different pruning granularities to adapt to varying initialization methods. In this paper, we evaluated the impact of four levels of granularities: block-wise, layer-wise, input-wise, and output-wise pruning, as shown in Table VI. For magnitude pruning, the block-wise granularity yields the best performance, while the output-wise granularity delivers the lowest perplexity for Wanda pruning. By tailoring the pruning granularity to the particular pruning approach, LLM-Barber can consistently achieve optimal model performance.
Figure 4: Ablation of calibration size in LLaMA3-8B. LLM-Barber is robust across varying calibration size.
TABLE VIII: Inference performance of different quantization methods on FPGA platform. Results are reported for Qwen2-7B on EdgeLLM with the ARC-Challenge, ARC-Easy, and BoolQ tasks in zero-shot setting.
Method ARC-Challenge ARC-Easy BoolQ
TPS $\uparrow$ Latency(ms) $\downarrow$ Acc(%) $\uparrow$ TPS $\uparrow$ Latency(ms) $\downarrow$ Acc(%) $\uparrow$ TPS $\uparrow$ Latency(ms) $\downarrow$ Acc(%) $\uparrow$
GPTQ 30.01 1.98 46.28 29.88 1.75 64.71 30.31 0.65 56.88
AWQ 28.13 2.23 35.38 28.85 2.02 57.62 28.47 1.98 51.84
APTQ+Ours 30.37 1.92 56.28 30.14 1.58 67.16 31.48 0.48 60.94
Pruning Metric. We analyze the effect of different pruning metrics, weight magnitude, gradient magnitude, and product of weight and gradient. As shown in Table VII, the product of weight and gradient consistently outperforms the others, achieving the lowest perplexity of 10.98 for Magnitude, 9.348 for SparseGPT, and 9.451 for Wanda. This confirms the effectiveness of our product-based pruning metric across various methods.
@@
IV-E Deployment-Aware Quantization Compatibility
TABLE IX: Comparative Throughput Analysis of Quantization Methods Across GPU Frameworks with LLM-Barber Optimization
Method Throughput(Tokens/sec) Baseline
CUDA Triton Marlin ExLlama BF16
GPTQ 37.38 15.90 38.03 50.59 44.30
AWQ 38.10 16.21 43.46 51.63 44.30
APTQ+Ours 38.29 16.42 47.13 52.29 44.30
LLM-Barber’s block-aware sparsity method provides an efficient compression solution when combined with post-training quantization like APTQ [12]. This integration enhances both model size reduction and inference speed. In this section, we present experimental results validating the compatibility and acceleration benefits of combining LLM-Barber with APTQ.
We evaluate the model’s performance on both GPU and FPGA platforms, comparing tokens per second (TPS), latency, and accuracy across different quantization methods: GPTQ [11], AWQ [30], and APTQ. As shown in Table IX, LLM-Barber combined with APTQ achieves higher TPS and lower latency on both platforms, while maintaining competitive accuracy.
Additionally, we extend this analysis to FPGA hardware using the EdgeLLM [31] platform, where we compare the performance of the quantization methods on the ARC-Challenge, ARC-Easy, and BoolQ datasets. As detailed in Table VIII, APTQ, when integrated with LLM-Barber’s block-sparsity, delivers the highest TPS, lowest latency, and best accuracy across all tasks. This validates the effectiveness of LLM-Barber’s block-aware sparsity and APTQ’s quantization on FPGA, demonstrating that our method significantly accelerates deployment while preserving model performance.
In summary, the combination of LLM-Barber’s block-sparsity approach with APTQ optimizes both inference speed and deployment flexibility. Our results confirm that this joint optimization provides higher TPS and lower latency on GPU while achieving substantial acceleration on FPGA platforms. These findings highlight the potential of integrating sparse pruning with quantization for end-to-end deployment of LLMs, enabling scalable, efficient, and hardware-agnostic solutions across different platforms, such as GPU and FPGA.
Future work will explore further quantization schemes and optimizations for edge deployment. We envision expanding this approach to larger model architectures and more diverse hardware environments, making large-scale LLM deployment more efficient and accessible.
V Conclusion
We introduced LLM-Barber, a one-shot pruning framework that achieves aggressive compression of large language models without retraining. By leveraging a block-aware mask rebuilding strategy and a weight-gradient importance metric, LLM-Barber removes 50–60% of weights while preserving state-of-the-art accuracy. Our approach prunes multi-billion-parameter models in minutes and incorporates hardware-aware optimizations, such as structured block sparsity and low-bit quantization, to ensure effective deployment on modern accelerators. Experimental results demonstrate that LLM-Barber outperforms unstructured pruning baselines like SparseGPT and Wanda in both perplexity and downstream task accuracy. Additionally, the FPGA deployment of a pruned+quantized model achieves significant speedups and energy savings, proving LLM-Barber’s hardware efficiency. In conclusion, LLM-Barber bridges the gap between theoretical model compression and practical deployment, enabling efficient LLM inference in resource-constrained environments. Future work will scale this approach to larger models and integrate it with other compression techniques.
Acknowledgement
This work was supported by the Shenzhen Science and Technology Program under Grant No. KQTD20200820113051096, Shenzhen Science and Technology Program under Grant No. JCYJ20220818100217038, the National Natural Science Foundation of China under Key Program Grant No. 62034007, the Science and Technology Innovation Key R&D Program of Chongqing under Grant No. CSTB2023TIAD-STX0001, and the Hong Kong Theme-based Research Scheme (TRS) project T45-701/22-R of the Research Grants Council (RGC).